# Biologically Plausible Learning via Bidirectional Spike-Based Distillation

### v1

@@ Abstract
Developing biologically plausible learning algorithms that can achieve performance comparable to error backpropagation remains a longstanding challenge. Existing approaches often compromise biological plausibility by entirely avoiding the use of spikes for error propagation or relying on both positive and negative learning signals, while the question of how spikes can represent negative values remains unresolved. To address these limitations, we introduce Bidirectional Spike-based Distillation (BSD), a novel learning algorithm that jointly trains a feedforward and a backward spiking network. We formulate learning as a transformation between two spiking representations (i.e., stimulus encoding and concept encoding) so that the feedforward network implements perception and decision-making by mapping stimuli to actions, while the backward network supports memory recall by reconstructing stimuli from concept representations. Extensive experiments on diverse benchmarks, including image recognition, image generation, and sequential regression, show that BSD achieves performance comparable to networks trained with classical error backpropagation. These findings represent a significant step toward biologically grounded, spike-driven learning in neural networks.
@@ 1 Introduction
We also show that the proposed bidirectional distillation (implemented via spiking trains) yields a more biologically plausible learning algorithm. While backpropagation has achieved remarkable success in deep learning (LeCun et al., 2015; Rumelhart et al., 1986), its underlying mechanisms conflict with established neurobiological principles (Crick, 1989; Lillicrap et al., 2020). Key inconsistencies include the requirement for symmetric feedforward and feedback weights, reliance on global error signals instead of local synaptic plasticity, a two-stage learning process that clearly separates forward and backward passes, and the use of continuous activations rather than discrete spike-based communication. To address these limitations, and building on the three criteria proposed by Lv et al. (2025), we introduce two additional requirements: neurons should communicate using discrete binary spikes for both learning and inference, and learning should rely on unsigned spiking signals only (Hayden et al., 2011). The learning algorithm we present demonstrates that all five criteria can be satisfied, whereas existing approaches typically fall short on one or more of these criteria.
@@ 2 Related Work
To address these limitations, numerous alternative approaches have been proposed to improve biological plausibility (Schmidgall et al., 2024; Jiao et al., 2022; Li et al., 2024). However, our method uniquely satisfies established biological constraints while maintaining strong performance and broad applicability across model architectures and tasks. For example, target propagation (TP) (Bengio, 2014) introduces approximate inverse models to generate layer-wise targets, with weight updates obtained by minimizing the mismatch between outputs and targets. Although TP employs local losses and avoids weight symmetry, it requires each layer to transmit two distinct types of signals at different times. Moreover, TP often suffers from convergence instability. Extensions such as Difference Target Propagation (DTP) (Lee et al., 2015) and SDTP (Bartunov et al., 2018) improved stability and performance but offered little progress towards greater biological plausibility, and Biologically-plausible Reward Propagation (BRP) (Zhang et al., 2021), which replaces floating-point interlayer signals in TP with spike-based signals, has not demonstrated reliable generalization across tasks. Other methods, such as Dendritic Localized Learning (DLL) (Lv et al., 2025), employ local error signals for weight updates, yet still rely on transmitting signed floating-point values and generally perform poorly across benchmarks. Counter-Current Learning (CCL) (Kao & Hariharan, 2024) faces the same limitation of floating-point communication and further lacks demonstrated effectiveness in sequential regression tasks.
@@
Other biologically inspired mechanisms, like Hebbian learning (Donald, 1949) and spike-timing-dependent plasticity (STDP) (Song et al., 2000), adjust synaptic strength according to correlations in neuronal activity, while burst-dependent synaptic plasticity(Payeur et al., 2021) regulates synaptic plasticity by high-frequency bursts of spikes. Although fully consistent with biological observations, these rules struggle to integrate supervised learning signals, and STDP additionally requires precise temporal resolution. Energy-based learning approaches (LeCun et al., 2006), e.g., Boltzmann machines (Ackley et al., 1985), Hopfield networks (Hopfield, 1984), and contrastive learning frameworks (Hinton, 2002), instead optimize an energy function. However, minimizing energy does not always correspond to reducing task-specific loss, limiting their utility for general supervised learning.
@@ 4.2 Model Architecture
To highlight the biological foundations of BSD, we first describe the neuron model employed in BSD. We follow Sacramento et al. (2018) and utilize pyramidal neurons with a three-compartment structure: soma, apical dendrites (carrying backward learning signals), and basal dendrites (receiving feedforward inputs). To satisfy criterion C4 (model of neurons), we use spiking neurons that emit discrete pulses instead of continuous activation values. For each neuron, the membrane potentials arriving at the soma from the basal and apical compartments are denoted $v$ and $\hat{v}$, respectively. The apical potential $\hat{v}$ acts as a supervisory signal guiding synaptic plasticity on the basal dendrites, while the basal potential $v$ drives spike generation through $s=\mathcal{SN}(v)$, where $\mathcal{SN}(\cdot)$ denotes the spiking operation that integrates input voltage and produces spike outputs.
@@
where $\mathbf{v}_{i}^{\prime}$ denotes the somatic membrane potential of Type 1 neurons in layer $i$ arising from basal dendritic integration, $\hat{\mathbf{v}}_{i}$ denotes the somatic membrane potential of Type 2 neurons in layer $i$ arising from apical dendritic integration. $\mathbf{W}_{i}$ is the synaptic weight for Type 1 neurons in layer $i$, and $\mathbf{s}_{i}$ denotes the spike train that Type 1 neurons in layer $i$ output.
@@ 4.3 Training Procedure
where $\lambda$ is a hyperparameter that controls the penalty strength for suppressing spurious correlations between non-corresponding voltage pairs. The local loss for Type 2 neurons, $\mathcal{L}^{\prime}_{i}$, is defined analogously, with its formulation provided in Appendix D. Compared with InfoNCE loss commonly used in contrastive learning scenarios, ReCo loss avoids penalizing negatively correlated voltage pairs between $\mathbf{V}_{i}$ and $\hat{\mathbf{V}}_{i}$, thereby introducing enhanced flexibility and representational richness to learned embeddings while preserving alignment objectives. Considering criterion C2 (local synaptic plasticity without global error signals), we employ detach() operations to truncate the computational graph between layers, ensuring that each neuron’s loss only propagates learning signals to synaptic weights connected to its dendrites. As a result, the gradient of the local loss $\mathcal{L}_{i}$ with respect to the feedforward weights $\mathbf{W}_{i-1}$ (for $i=2,\dots,L-1$) is given by:
@@
where $B$ denotes the batch size, $\mathbf{s}_{i-1,k}$ represents the spike output vector of Type 1 neurons from layer i1i-1 for the $k$-th sample, and $\mathcal{L}_{i}$ denotes the local loss for neurons in layer $i$. The gradient update formula for $\bm{\Theta}$ takes an analogous form, and the detailed gradient derivations for both $\mathbf{W}$ and $\mathbf{\Theta}$ are presented in Appendix D. The total loss for Type 1 neurons is defined as:
@@ 4.4 Learning for generation tasks
We employ an autoencoder architecture for generation tasks, where both bottom-up and top-down inputs are images $\mathbf{x}$. To capture fine-grained edge details while reducing noise artifacts, we apply Fast Fourier Transform (FFT) decomposition to the input images as well as to all basal and apical membrane voltages. This frequency-domain representation separates low- and high-frequency components, enabling adaptive loss computation. Within each layer, the regularization parameter $\lambda$ is adjusted according to frequency content: larger values are assigned to high-frequency components to suppress spurious correlations and preserve edge fidelity, whereas smaller values are used for low-frequency components to avoid noise amplification and maintain structural coherence. At the top layer, mean squared error is employed to compute the loss between basal and apical voltages.
@@ 5.2 Implementation Details
To ensure fair comparison, identical network architectures are used across all learning algorithms for each task. For MLPs, CNNs, RNNs, and autoencoders, the same configurations are applied when comparing methods on a given dataset. Detailed dataset descriptions, architectural specifications, hyperparameter settings, and evaluation metrics are provided in Appendix E.
@@
Backpropagation
on ANNs
MLPs $\bf 98.77\%$$\pm 0.33\%$ $\bf 89.59\%$$\pm 0.14\%$ $\bf 61.65\%$$\pm 0.42\%$ $\bf 57.65\%$$\pm 0.08\%$ $\bf 27.92\%$$\pm 0.17\%$ $\bf 67.12\%$
@@
Backpropagation
on SNNs
MLPs $98.57\%$$\pm 0.24\%$ $88.87\%$$\pm 0.65\%$ $61.28\%$$\pm 0.52\%$ $49.95\%$$\pm 0.53\%$ $23.32\%$$\pm 0.87\%$ $64.40\%$
@@ 5.3 Image Classification
Our proposed BSD algorithm reconciles biological plausibility with competitive performance. BSD satisfies all five criteria of biological plausibility (C1C5) while achieving performance comparable to backpropagation, with stable convergence across diverse datasets. In particular, on challenging datasets such as SVHN, CIFAR-10, and CIFAR-100, and on more complex architectures including CNNs, BSD consistently delivers robust performance. Taken together, these findings underscore the feasibility of attaining biological plausibility without sacrificing task performance.
@@ 5.6 Ablation Study
As mentioned in Section 4, inspired by contrastive learning, we adopt ReCo loss as our layer-wise loss function. Here, we delve deeper into the impact of layer-wise loss function selection, analyzing how BSD performs when using MSE or InfoNCE (van den Oord et al., 2018), a loss commonly used in contrastive learning scenarios, for aligning intraneuronal voltages. We term the method using MSE as layer-wise loss ”BSD-MSE” and the method using InfoNCE as ”BSD-InfoNCE,” and conduct ablation experiments on both image classification and sequential regression tasks.
@@
To conclude, employing MSE loss for intraneuronal voltage alignment leads to convergence difficulties, indicating that it is ill-suited for BSD training. Moreover, compared to InfoNCE, ReCo loss offers a distinct advantage by not penalizing unpaired samples that are already orthogonal or negatively correlated, which allows for more flexible alignment and richer feature representations.
@@ 5.7 Training Analysis
In this section, we analyze the convergence behavior of BSD-trained models and the properties of Type 1 and Type 2 neurons during training. Since BSD applies ReCo loss to all layers except the top one and relies on spikes for inter-neuronal communication, both the penalty weight $\lambda$ in ReCo loss and the neuronal firing threshold are critical factors influencing its convergence.
Figure 3(a) shows the effect of the penalty weight $\lambda$ on BSD-trained CNNs evaluated on the SVHN dataset. Performance drops markedly when $\lambda=0$, underscoring the role of $\lambda$ in promoting separation among sample representations and thereby enlarging the representational space. We further observe that smaller values of $\lambda$ accelerate convergence, while the best final performance is achieved at $\lambda=0.6$. Figure 3(b) illustrates the relationship among neuronal firing threshold, spike sparsity, and network performance. The network attains its best performance at a firing threshold of 0.6, suggesting that thresholds that are either too low or too high impair effective learning. Figure 3(c) shows the Hamming similarity between the spike trains of Type 1 and Type 2 neurons within the same layer. The similarity rises quickly during training and exceeds 0.85 by epoch 20, demonstrating that the feedforward (stimuli-to-decision) and backward (concept-to-stimuli) pathways successfully achieve mutual alignment of their spiking feature representations through bidirectional distillation, thereby validating the effectiveness of our joint training framework.
For completeness, we provide $t$-SNE visualizations of representations on BSD-trained CNNs in Appendix F.5 and examine the degree of alignment between weights $\mathbf{W}$ and $\bm{\Theta}$ in Appendix F.4.

### v2

@@ Abstract
Developing biologically plausible learning algorithms that can achieve performance comparable to error backpropagation remains a longstanding challenge. Existing approaches often compromise biological plausibility by entirely avoiding the use of spikes for error propagation or relying on both positive and negative learning signals, while the question of how spikes can represent negative values remains unresolved. To address these limitations, we introduce Bidirectional Spike-based Distillation (BSD), a novel learning algorithm that jointly trains a feedforward and a backward spiking network. We formulate learning as a transformation between two spiking representations (i.e., stimulus encoding and concept encoding) so that the feedforward network implements perception and decision-making by mapping stimuli to actions, while the backward network supports memory recall by reconstructing stimuli from concept representations. Extensive experiments on diverse benchmarks, including image recognition, image generation, and sequential regression, show that BSD achieves performance comparable to networks trained with classical error backpropagation. These findings represent a significant step toward biologically grounded, spike-driven learning in neural networks. Our code is available at https://github.com/alden199/Bidirectional-Spike-Based-Distillation.
@@ 1 Introduction
We also show that the proposed bidirectional distillation (implemented via spike trains) yields a more biologically plausible learning algorithm. While backpropagation has achieved remarkable success in deep learning (LeCun et al., 2015; Rumelhart et al., 1986), its underlying mechanisms conflict with established neurobiological principles (Crick, 1989; Lillicrap et al., 2020). Key inconsistencies include the requirement for symmetric feedforward and feedback weights, reliance on global error signals instead of local synaptic plasticity, a two-stage learning process that clearly separates forward and backward passes, and the use of continuous activations rather than discrete spike-based communication. To address these limitations, and building on the three criteria proposed by Lv et al. (2025), we introduce two additional requirements: neurons should communicate using discrete binary spikes for both learning and inference, and learning should rely on unsigned spiking signals only (Hayden et al., 2011). The learning algorithm we present demonstrates that all five criteria can be satisfied, whereas existing approaches typically fall short on one or more of these criteria.
@@ 2 Related Work
To address these limitations, numerous alternative approaches have been proposed to improve biological plausibility (Schmidgall et al., 2024; Jiao and others, 2022; Li and others, 2024). Recent advances in spiking neural networks have introduced attention mechanisms (Yao et al., 2023b; a) and residual learning (Hu et al., 2024) to enhance representational capacity while maintaining spike-based communication. To eliminate the reliance on global error signals, local loss methods (Mostafa et al., 2018) and their variants (Belilovsky et al., 2019; Nøkland and Eidnes, 2019; Kaiser et al., 2020) have been introduced, which typically employ fixed or trainable auxiliary heads to align hidden layers directly with the target. Feedback alignment (Lillicrap et al., 2016) addresses the weight transport problem by replacing the backward weights with fixed, randomly initialized matrices, thereby breaking the symmetry between forward and backward weights. Target propagation (TP) (Bengio, 2014) introduces approximate inverse models to generate layer-wise targets, with weight updates obtained by minimizing the mismatch between outputs and targets. Although TP employs local losses and avoids weight symmetry, it requires each layer to transmit two distinct types of signals at different times. Moreover, TP often suffers from convergence instability. Extensions such as Difference Target Propagation (DTP) (Lee et al., 2015) and SDTP (Bartunov et al., 2018) improved stability and performance but offered little progress towards greater biological plausibility, and Biologically-plausible Reward Propagation (BRP) (Zhang et al., 2021), which replaces floating-point interlayer signals in TP with spike-based signals, has not demonstrated reliable generalization across tasks. Predictive coding (Rao and Ballard, 1999) offers another influential framework, postulating that the brain continually generates top-down predictions to minimize sensory prediction errors. Motivated by this framework, Error-driven Local Representation Alignment (LRA-E) (Ororbia and Mali, 2019) introduces a mechanism where error signals are projected backward to generate local target representations for hidden layers. The network then learns by minimizing the local discrepancy between actual neuronal activities and these generated targets, thereby enabling training without global error backpropagation. Other methods, such as Decoupled Neural Interfaces (DNI) (Jaderberg et al., 2017), train an auxiliary head at each hidden layer to approximate layer-wise gradients yet fail to break weight symmetry. Alternatively, Dendritic Localized Learning (DLL) (Lv et al., 2025) employs local error signals for weight updates, yet still relies on transmitting signed floating-point values and generally performs poorly across benchmarks. Counter-Current Learning (CCL) (Kao and Hariharan, 2024) faces the same limitation of floating-point communication and further lacks demonstrated effectiveness in sequential regression tasks.
@@
Other biologically inspired mechanisms, like Hebbian learning (Donald, 1949) and spike-timing-dependent plasticity (STDP) (Song et al., 2000), adjust synaptic strength according to correlations in neuronal activity, while burst-dependent synaptic plasticity(Payeur et al., 2021) regulates synaptic plasticity by high-frequency bursts of spikes. Although fully consistent with biological observations, these rules struggle to integrate supervised learning signals, and STDP additionally requires precise temporal resolution. Energy-based learning approaches (LeCun et al., 2006), e.g., Boltzmann machines (Ackley et al., 1985), Hopfield networks (Hopfield, 1984), and contrastive learning frameworks (Hinton, 2002), instead optimize an energy function. However, minimizing energy does not always correspond to reducing task-specific loss, limiting their utility for general supervised learning. E-prop (Bellec et al., 2020) approximates BPTT for SRNNs using eligibility traces; while supporting local, online updates, it remains limited by the use of signed error signals.
@@ 4.2 Model Architecture
To highlight the biological foundations of BSD, we first describe the neuron model employed in BSD. We follow Sacramento et al. (2018) and utilize pyramidal neurons with a three-compartment structure: soma, apical dendrites (carrying backward learning signals), and basal dendrites (receiving feedforward inputs). To satisfy criterion C4 (model of neurons), we use spiking neurons that emit discrete pulses instead of continuous activation values. For each neuron, the membrane potentials arriving at the soma from the basal and apical compartments are denoted $v$ and $\hat{v}$, respectively. The apical potential $\hat{v}$ acts as a supervisory signal guiding synaptic plasticity on the basal dendrites, while the basal potential $v$ drives spike generation through $s=\mathcal{SN}(v)$, where $\mathcal{SN}(\cdot)$ denotes the spiking operation that integrates input voltage and produces spike outputs. In the following sections, we use bold lowercase letters (e.g., $\mathbf{v}_{i}$, $\hat{\mathbf{v}}_{i}$) to denote layer-wise vectors, which are formed by aggregating the single-neuron scalar potentials ($v$ and $\hat{v}$, respectively) from all neurons in layer $i$.
@@
where $\mathbf{v}_{i}$ denotes the somatic membrane potential of Type 1 neurons in layer $i$ from basal dendritic integration, and $\hat{\mathbf{v}}_{i}^{\prime}$ denotes the somatic membrane potential of Type 2 neurons in layer $i$ arising from apical dendritic integration. $\mathbf{W}_{i}$ is the synaptic weight for Type 1 neurons in layer $i$, and $\mathbf{s}_{i}$ denotes the spike train that Type 1 neurons in layer $i$ output.
@@ 4.3 Training Procedure
where $\lambda$ is a hyperparameter that controls the penalty strength for suppressing spurious correlations between non-corresponding voltage pairs. The local loss for Type 2 neurons, $\mathcal{L}^{\prime}_{i}$, is defined analogously, with its formulation provided in Appendix D. Compared with InfoNCE loss commonly used in contrastive learning scenarios, ReCo loss avoids penalizing negatively correlated voltage pairs between $\mathbf{V}_{i}$ and $\hat{\mathbf{V}}_{i}$, thereby introducing enhanced flexibility and representational richness to learned embeddings while preserving alignment objectives. We provide a more detailed justification for adopting the ReCo loss in Appendix Q. Considering criterion C2 (local synaptic plasticity without global error signals), we employ detach() operations to truncate the computational graph between layers, ensuring that each neuron’s loss only propagates learning signals to synaptic weights connected to its dendrites. As a result, the gradient of the local loss $\mathcal{L}_{i}$ with respect to the feedforward weights $\mathbf{W}_{i-1}$ (for $i=2,\dots,L-1$) is given by:
@@
where $B$ denotes the batch size, $\mathbf{s}_{i-1,k}$ represents the spike output vector of Type 1 neurons from layer i1i-1 for the $k$-th sample, and $\mathcal{L}_{i}$ denotes the local loss for neurons in layer $i$. This locally computed gradient is then used to update the feedforward synaptic weights via a standard gradient descent step for $i=2,\dots,L-1$:
$\mathbf{W}_{i-1}^{\text{new}}=\mathbf{W}_{i-1}^{\text{old}}+\eta_{\mathbf{W}}\frac{\partial\mathcal{L}_{i}}{\partial\mathbf{W}_{i-1}},$ (9)
where the superscripts ’new’ and ’old’ denote the weights after and before the update, and $\eta_{\mathbf{W}}$ is the learning rate for the feedforward weights. A symmetric update rule, also based on its corresponding local gradient, is applied to the backward weights $\bm{\Theta}$. The detailed gradient derivations for both $\mathbf{W}$ and $\mathbf{\Theta}$ are presented in Appendix D. The total loss for Type 1 neurons is defined as:
@@ 4.4 Learning for generation tasks
We employ an autoencoder architecture for generation tasks, where both bottom-up and top-down inputs are images $\mathbf{x}$. To capture fine-grained edge details while reducing noise artifacts, we apply Fast Fourier Transform (FFT) decomposition to the input images as well as to all basal and apical membrane voltages. This frequency-domain representation separates low- and high-frequency components, enabling adaptive loss computation. Within each layer, the regularization parameter $\lambda$ is adjusted according to frequency content: larger values are assigned to high-frequency components to suppress spurious correlations and preserve edge fidelity, whereas smaller values are used for low-frequency components to avoid noise amplification and maintain structural coherence. At the top layer, mean squared error is employed to compute the loss between basal and apical voltages. The detailed mechanism of the FFT decomposition is explained in Appendix M.
@@ 5.2 Implementation Details
To ensure fair comparison, identical network architectures are used across all learning algorithms for each task. For MLPs, CNNs, RNNs, and autoencoders, the same configurations are applied when comparing methods on a given dataset. Detailed dataset descriptions, architectural specifications, hyperparameter settings, and evaluation metrics are provided in Appendix E. A comprehensive analysis of the computational costs, including training and inference memory consumption, is provided in Appendix N.
@@
Backpropagation on ANNs MLPs $\bf 98.77\%$$\pm 0.33\%$ $\bf 89.59\%$$\pm 0.14\%$ $\bf 61.65\%$$\pm 0.42\%$ $\bf 57.65\%$$\pm 0.08\%$ $\bf 27.92\%$$\pm 0.17\%$ $\bf 67.12\%$
@@
Backpropagation on SNNs MLPs $98.57\%$$\pm 0.24\%$ $88.87\%$$\pm 0.65\%$ $61.28\%$$\pm 0.52\%$ $49.95\%$$\pm 0.53\%$ $23.32\%$$\pm 0.87\%$ $64.40\%$
@@ 5.3 Image Classification
Our proposed BSD algorithm reconciles biological plausibility with competitive performance. BSD satisfies all five criteria of biological plausibility (C1C5) while achieving performance comparable to backpropagation, with stable convergence across diverse datasets. In particular, on challenging datasets such as SVHN, CIFAR-10, and CIFAR-100, and on more complex architectures including CNNs, BSD consistently delivers robust performance. Taken together, these findings underscore the feasibility of attaining biological plausibility without sacrificing task performance. To further assess the scalability of our approach, we conducted experiments on the more complex Tiny-ImageNet dataset, with the results presented in Appendix I. We also evaluate the robustness of our models against input noise in Appendix L.
@@ 5.6 Ablation Study
As mentioned in Section 4, inspired by contrastive learning, we adopt ReCo loss as our layer-wise loss function. Here, we delve deeper into the impact of layer-wise loss function selection, analyzing how BSD performs when using MSE or InfoNCE (van den Oord et al., 2018), a loss commonly used in contrastive learning scenarios, for aligning intraneuronal voltages. We term the method using MSE as layer-wise loss “BSD-MSE” and the method using InfoNCE as “BSD-InfoNCE,” and conduct ablation experiments on both image classification and sequential regression tasks.
@@
To conclude, employing MSE loss for intraneuronal voltage alignment leads to convergence difficulties, indicating that it is ill-suited for BSD training. Moreover, compared to InfoNCE, ReCo loss offers a distinct advantage by not penalizing unpaired samples that are already orthogonal or negatively correlated, which allows for more flexible alignment and richer feature representations. We also investigate the impact of the number of timesteps on model performance in Appendix O. An ablation study demonstrating the importance of Batch Normalization is presented in Appendix P.
@@ 5.7 Training Analysis
In this section, we analyze the convergence behavior of BSD-trained models and the properties of Type 1 and Type 2 neurons during training. Since BSD applies ReCo loss to all layers except the top one and relies on spikes for inter-neuronal communication, both the penalty weight $\lambda$ in ReCo loss and the neuronal firing threshold are critical factors influencing its convergence. Additionally, we investigate the sensitivity of our framework to batch size in Appendix J, as contrastive learning methods often depend on sufficiently large batches. A detailed analysis of the energy efficiency of our BSD-trained models during inference is provided in Appendix K, highlighting the benefits of spike-based computation.
Figure 3(a) shows the effect of the penalty weight $\lambda$ on BSD-trained CNNs evaluated on the SVHN dataset. Performance drops markedly when $\lambda=0$, underscoring the role of $\lambda$ in promoting separation among sample representations and thereby enlarging the representational space. We further observe that smaller values of $\lambda$ accelerate convergence, while the best final performance is achieved at $\lambda=0.6$. Figure 3(b) illustrates the relationship among neuronal firing threshold, spike sparsity, and network performance on SVHN. The network attains its best performance at a firing threshold of 0.6, suggesting that thresholds that are either too low or too high impair effective learning. Figure 3(c) shows the Hamming similarity between the spike trains of Type 1 and Type 2 neurons within the same layer. The similarity rises quickly during training and exceeds 0.85 by epoch 20, demonstrating that the feedforward (stimuli-to-decision) and backward (concept-to-stimuli) pathways successfully achieve mutual alignment of their spiking feature representations through bidirectional distillation, thereby validating the effectiveness of our joint training framework.
For completeness, we provide $t$-SNE visualizations of representations on BSD-trained CNNs in Appendix H and examine the degree of alignment between weights $\mathbf{W}$ and $\bm{\Theta}$ in Appendix G.